A CURRENT OF 9.65 AMPERE IS PASSED THROUGH THE AQUEOUS SOLUTION Nacl  USING SUITABLE ELECTRODES FOR 1000s. THE AMOUNT OF NaOH FOEMED DURING ELECTROLYSIS IS 

Grade:12th pass

1 Answers

Arun
4 years ago
The liberation of each h+ ion can result in 1 naoh being formed .moles of electron formed = (9.65×1000)÷96500 = 0.1so moles of h+ liberated =0.1 moles of naoh = 0.1 mass of naoh formed = 40×0.1= 4.0 g